OpenStudy (anonymous):

The length of a rectangular garden is 5 m greater than the width. The area is 84 m^2. Find the dimensions of the garden.

OpenStudy (anonymous):

let the length be x, then breadth=x-5 area =L*B x(x-5)=84 x^2-5x-84=0 (x+7)(x-12)=0 x=-7 or x=12 take the positive value IE 12m as the answer for length breadth = x-5=12-5=7m